### v1

# Families of cyclic curve coverings with maximal monodromy

Abstract.
We study the algebraic monodromy of families of cyclic Galois coverings of curves. Under a condition on the $G$-decomposition of the associated variation of Hodge structures, we prove a criterion for the maximality of the monodromy. The proof combines the genus-zero case with a degeneration argument involving Prym varieties of certain admissible coverings. As a consequence of our criterion, we show that for $g\geq 8$ there exists no special family of Galois covers of the type we consider, providing new evidence towards the Coleman-Oort conjecture. Finally, we determine when the loci of double and triple Galois covers are totally geodesic.
2020 Mathematics Subject Classification. 14H10, 14H40, 14G35, 14D07.
The authors are members of GNSAGA (INdAM). I. Spelta was partially supported by the Spanish MINECO research project PID2023-147642NB-I00. C.Tamborini was partially supported by the DFG-Research Training Group 2553 “Symmetries and classifying spaces: analytic, arithmetic, and derived”
1. Introduction
$1.$
In this paper, we prove a criterion for maximality of the algebraic monodromy group for certain families of cyclic coverings of curves. Our technique relies on the study of the natural $\mathbb{Q}$}-variation of Hodge structures associated with the family. For a cyclic group $G$, let $f:\mathcal{C}\rightarrow B$ be the family parametrizing all Galois covers $C_{b}\to C^{\prime}_{b}=C_{b}/G$ with prescribed ramification and monodromy. The $G$-action on the curves induces a $G$-action on $H^{1}(C_{b},{\mathbb{Q}})$, which preserves the cup product and yields the inclusion $G\subset\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))$. Let $\operatorname{Mon}^{0}\subset\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))$ be the algebraic monodromy group associated with the semisimple local system $\mathbb{V}=R^{1}f_{*}{\mathbb{Q}}$. Since the monodromy must preserve the cup product as well as the $G$-action, we have the inclusion
$\operatorname{Mon}^{0}\;\subseteq\;(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der},$
where $\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G}$ is the centralizer of $G$ inside $\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))$, and $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}$ denotes its derived subgroup. It is natural to investigate when such inclusion is actually an equality. To address this problem, we exploit the following decomposition induced by the group $G$:
(1.1) $\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G}\simeq\prod_{(\chi,\bar{\chi}),\chi\neq\bar{\chi}}\operatorname{U}(m_{\chi},m_{\bar{\chi}})\times\prod_{(\chi,\bar{\chi}),\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}),$
where $m_{\chi}$ is the multiplicity of the $\chi$-eigenspace for the action of $G$ on $H^{1,0}(C_{b})$ (and similarly for the conjugate $\bar{\chi}$). Taking the adjoint in (1.1), one forgets the factors coming from 0-dimensional eigenspaces. In the following, we consider families with no repeating factors in the monodromy, meaning that, for any two characters $\chi_{1}$ and $\chi_{2}$ with $m_{\chi_{i}},m_{\bar{\chi}_{i}}\neq 0$, the multiplicities satisfy $\{m_{\chi_{1}},\,m_{\bar{\chi}_{1}}\}\neq\{m_{\chi_{2}},\,m_{\bar{\chi}_{2}}\}$. In other words, a family has no repeating factors in the monodromy if all the factors appearing in $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}$ are mutually distinct. Our first result is the following.
Theorem A.
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then the monodromy group is maximal, namely
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}.$
In particular, if $g^{\prime}=g(C^{\prime})\geq 2$, we have
$\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}.$
(See Theorem 4.7). Maximality results for the monodromy of families of algebraic varieties are widely studied in algebraic geometry. Regarding $G$-coverings of curves, we mention e.g. [4, 3, 19, 17]. In particular we point out that, for $g^{\prime}\geq 4$, Theorem A partially recovers [17, Theorem 1.3], where the authors prove a strong maximality result for the monodromy of families of any Galois covers, requiring some lower bound on the genus of the base curve $C^{\prime}$. By studying the adjoint monodromy group $\operatorname{Mon}^{0,ad}$, we ignore the contribution of 0-dimensional eigenspaces and, thus, we can take into account lower genus $g^{\prime}$, too. In fact, our proof proceeds by induction on the genus $g^{\prime}$. The case $g^{\prime}=0$ follows from [26]. The induction step is based on a degeneration argument and requires the study of Prym varieties associated with certain admissible coverings, which is carried out in Theorem 3.3. As a consequence, under the assumption of cyclic coverings with no repeating factors, our argument applies to all genera $g^{\prime}$, and in the case $g^{\prime}\geq 4$ it provides an alternative proof of the maximality of the monodromy. Notice that, if $g^{\prime}\geq 2$, no 0-dimensional eigenspace appears and, hence, our result concerns the full monodromy group $\operatorname{Mon}^{0}$.
1.2.
The maximality of the monodromy is a powerful property and it has strong consequences on the geometry of the moduli image of our families. In particular, our result has interesting applications to the Coleman-Oort conjecture. Denote by $\mathcal{M}_{g}$ the moduli space of smooth projective curves of genus $g$ over $\mathbb{C}$, by $\mathcal{A}_{g}$ the moduli space of $g$-dimensional principally polarized abelian varieties, and let $j:\mathcal{M}_{g}\to\mathcal{A}_{g}$ be Torelli map sending curves to their Jacobians. The Coleman-Oort conjecture concerns the interaction between the image of the Torelli map and the special subvarieties of $\mathcal{A}_{g}$, which are defined as Hodge loci associated with the natural rational variation of Hodge structure on $\mathcal{A}_{g}$. Special subvarieties are totally geodesic, namely they are linear with respect to the locally symmetric metric induced from the Siegel space. The conjecture predicts that for $g\gg 0$, there are no positive-dimensional special subvarieties $S\subset\mathcal{A}_{g}$ such that $S$ is generically contained in the Torelli locus, i.e. $S\subset\overline{j(\mathcal{M}_{g})}$ and $S\cap j(\mathcal{M}_{g})\neq\emptyset$. However, for low genus, counterexamples are known: there exist positive dimensional special subvarieties $S\subset\mathcal{A}_{g}$ that lie generically in the Torelli locus. All known examples occur for $g\leq 7$, and arise from families of Galois covers of curves. Several such examples have been constructed and studied in the literature; see for instance [8, 24, 22, 23, 10, 11, 13, 16, 20, 25, 27]; we recall their main construction in section 5. The classification of such examples in case of cyclic covering of $\mathbb{P}^{1}$ is due to Moonen ([22]). The complete list of all currently known examples of special families of Galois coverings is provided in [10, 11]. Our second result provides new evidence for the Coleman-Oort conjecture.
Theorem B.
For $g\geq 8$ there are no totally geodesic (in particular, no special) families of cyclic coverings with no repeating factors in the monodromy.
(See Theorem 5.5). More precisely, we prove that the only positive dimensional totally geodesic families of cyclic coverings with no repeating factors in the monodromy are among the known ones (namely, among the ones appearing in [10, 11]). Moreover, we observe that, even though assuming no repeating factors appears as a strong condition on the monodromy, we have:
Proposition.
All known examples of special families of cyclic coverings (namely, the ones appearing in [10, 11]) have no repeating factors in the monodromy.
(See Theorem 5.8). Theorem B follows combining Theorem A with the fact that the smallest totally geodesic subvariety containing the moduli image of a family is given by (the image in $\mathcal{A}_{g}$) of the orbit of $\operatorname{Mon}^{0,ad}_{\mathbb{R}}$. When $g^{\prime}=0$, our results recover Moonen’s classification of families of cyclic covers of the projective line under the assumption that the monodromy has no repeating factors [22]. Furthermore, this assumption allows us to treat the case $g^{\prime}>0$ as well. We point out that, along with the proof, we also show that families with no repeating factors in the monodromy are special if and only if they are totally geodesic (see Proposition 5.4).
1.3.
Simple examples of families with no repeating factors in the monodromy are provided by families of double and triple coverings. Let $\mathcal{B}_{g,h}$ be the image in $\mathcal{M}_{g}$ of the family of all double covers $f:C\rightarrow C^{\prime}$ with $g(C)=g$ and $g(C^{\prime})=h$. Likewise, let $\mathcal{T}_{g,h}(r,m)$ be the image in $\mathcal{M}_{g}$ of the family of all Galois triple coverings $f:C\rightarrow C^{\prime}$ with $g(C)=g,\ g(C^{\prime})=h$, $r=2m+n$ ramification points and fixed monodromy datum $\Theta=(x_{1},\ldots,x_{2m},y_{1},\ldots,y_{n})$, where $x_{i}+x_{i+1}=0$ mod $3$ and $y_{i}+y_{j}\neq 0$ mod $3$ for all $i$ and $j$. Our third result is:
Theorem C.
The double covering locus $\mathcal{B}_{g,h}$ is not totally geodesic unless
$(g,h)=(1,0),(2,0),(2,1),(3,1).$
The triple covering locus $\mathcal{T}_{g,h}(r,m)$ is not totally geodesic unless
(g,h,r,m)=(2,0,4,2),(3,0,5,1),(4,0,6,0),(3,1,2,1),(4,1,3,0).(g,h,r,m)=(2,0,4,2),(3,0,5,1),(4,0,6,0),(3,1,2,1),(4,1,3,0).
(See Corollaries 6.1, 6.2, 6.3). The case of the bielliptic locus $\mathcal{B}_{g,1}$ was already considered in [12], where, with a completely different technique, the authors show that it is not totally geodesic for $g\geq 4$ (while it is for $g=3$). Our Theorem $C$ extends this result to all double and triple covering loci.
The paper is organized as follows: section 2 recalls some basic facts on families of coverings, section 3 discusses a result on Prym varieties of certain admissible coverings, section 4 proves Theorem A, while Theorem B is shown in section 5. Finally, in section 6 we study families of double and triple coverings.
Acknowledgements. We thank Matteo Costantini, Juan Carlos Naranjo, and Andrés Rojas for useful comments.
2. Preliminaries on families of $G$-curves
Let $G$ be a finite group and let $C,\widetilde{C}$ be smooth projective complex curves equipped with a $G$-action. We say that $C$ and $\widetilde{C}$ have the same topological type (or are topologically equivalent) if there exists an automorphism $\eta\in\operatorname{Aut}(G)$ together with an orientation–preserving homeomorphism
$f:C\to\widetilde{C}$
such that
$f(g\cdot x)=\eta(g)\cdot f(x),\qquad g\in G,\;x\in C.$
If, in addition, the map $f$ is biholomorphic, then $C$ and $\widetilde{C}$ are called $G$-isomorphic.
We denote by $\mathfrak{T}_{g^{\prime},r}(G)$ the collection of topological types of $G$-actions for which the quotient $C^{\prime}:=C/G$ has genus $g^{\prime}$ and the covering $C\to C^{\prime}$ has exactly $r$ branch points. Given a fixed topological type $\tau\in\mathfrak{T}_{g^{\prime},r}(G)$, there exists an algebraic family of curves with a $G$-action (see [15, 14]),
$\pi:\mathcal{C}_{\tau}\longrightarrow\mathsf{B}_{\tau},$
with the following properties:
  1. (1)
    every fiber in the family carries a $G$-action topologically equivalent to $\tau$;
  2. (2)
    any curve endowed with a $G$-action of type $\tau$ is $G$-isomorphic to some fiber, and only finitely many fibers represent the same isomorphism class.
In the following, we denote by $\mathsf{M}_{\tau}$ the image of the family in $\mathcal{M}_{g}$ and by $\mathsf{Z}_{\tau}$ the closure of its image in $\mathcal{A}_{g}$ under the Torelli map.
It is a classical fact that topological types of $G$-actions on curves can be described combinatorially in terms of monodromy and branching data. More concretely, fix a compact Riemann surface $C^{\prime}$ of genus $g^{\prime}\geq 0$ and an ordered $r$-tuple of distinct points $t=(t_{1},\dots,t_{r})$ on $C^{\prime}$. Let $U_{t}:=C^{\prime}-\{t_{1},\dots,t_{r}\}$ and choose a base point $t_{0}\in U_{t}$. Then the fundamental group $\pi_{1}(U_{t},t_{0})$ admits a canonical presentation
$\pi_{1}(U_{t},t_{0})\;\cong\;\Gamma_{g^{\prime},r}\;=\;\big\langle\alpha_{1},\beta_{1},\dots,\alpha_{g^{\prime}},\beta_{g^{\prime}},\gamma_{1},\dots,\gamma_{r}\;\big|\;\prod_{i=1}^{r}\gamma_{i}\cdot\prod_{j=1}^{g^{\prime}}[\alpha_{j},\beta_{j}]=1\big\rangle,$
where $\alpha_{j},\beta_{j}$ are simple loops generating a symplectic basis of $H_{1}(C^{\prime},\mathbb{Z})$, and $\gamma_{i}$ is a small loop around the puncture $t_{i}$, defined with respect to a chosen system of arcs $\tilde{\gamma}_{i}$ from $t_{0}$ to $t_{i}$.
Now suppose $f:C\to C^{\prime}$ is a $G$-Galois covering branched over $t$. Writing $V=f^{-1}(U_{t})$, the restriction $f|_{V}:V\to U_{t}$ is an unramified Galois covering with deck transformation group $G$. Hence there is a surjective homomorphism
$\pi_{1}(U_{t},t_{0})\twoheadrightarrow G,$
well-defined up to inner automorphisms of $G$. Equivalently, we obtain an epimorphism
$\theta:\Gamma_{g^{\prime},r}\twoheadrightarrow G.$
If $m_{i}$ denotes the order of the local monodromy around $t_{i}$, we set $m=(m_{1},\dots,m_{r})$, and we recall the following notion.
Definition 2.1.
A datum is a triple $(m,G,\theta)$, where
  • $m=(m_{1},\dots,m_{r})$ is an $r$-tuple of integers $m_{i}\geq 2$;
  • $G$ is a finite group;
  • $\theta:\Gamma_{g^{\prime},r}\twoheadrightarrow G$ is an epimorphism such that $\theta_{i}:=\theta(\gamma_{i})$ has order $m_{i}$ for each $i$.
The topological type of a $G$-action on a curve is thus encoded by such a datum: the group $G$, the local monodromy orders $m_{i}$, and the epimorphism $\theta$. As a consequence, the classification of topological types reduces to the combinatorics of generating systems of $G$ subject to these relations.
Finally, let us recall that for any fiber $C_{b},\ b\in\mathsf{B}_{\tau}$ of the family, the group action induces an action of $G$ on holomorphic one-forms:
$\rho:G\rightarrow\operatorname{GL}(H^{0}(C_{b},K_{C_{b}})),\;\rho(g)(\omega)=g.\omega:=(g^{-1})_{*}(\omega).$
Notice that the equivalence class of $\rho$ does not depend on the point $b\in B_{\tau}$. The Chevalley-Weil formula computes the dimension $m_{\chi}$ of the $\chi$-eigenspace for the action of $G$ on $H^{0}(C,K_{C})$, where $\chi$ is a $\mathbb{C}$-irreducible representation of $G$. In particular, for a family of cyclic coverings of order $d$, with monodromy $\Theta=(\theta_{1},\ldots,\theta_{r})$ and genus of the base curve $g^{\prime}$, the formula reads as follows ([6]):
(2.1) $m_{\chi_{i}}=(g^{\prime}-1)+\sum_{i=1}^{r}\frac{[i\theta_{i}]_{d}}{d},$
where $\chi_{i}$ is the representation associated with the eigenvalue $\xi^{i}$, being $\xi$ a primitive $d$-th root of the unity, and $a]_{d}$ denotes the unique representative of $a\in\mathbb{Z}/d\mathbb{Z}$ in $\{0,...,d-1\}$.
3. An admissible covering and its Prym variety
In this section, we consider admissible $G$-coverings of connected stable curves of arithmetic genus $g^{\prime}$ and we investigate their associated Prym varieties. We start by recalling the definition.
Definition 3.1.
Let $\mathcal{X}\rightarrow S$ be a family of connected nodal curves of arithmetic genus g and let $G$ be a finite group. A $G$-covering $f:\mathcal{Z}\rightarrow\mathcal{X}$ is called a family of admissible $G$-coverings if it satisfies
  • 1)
    the composition $\mathcal{Z}\rightarrow\mathcal{X}$ is a family of nodal curves;
  • 2)
    every node of a fiber of $\mathcal{Z}\rightarrow S$ maps to a node of the corresponding fiber of $\mathcal{X}\rightarrow S$;
  • 3)
    $\mathcal{Z}\rightarrow\mathcal{X}$ is a principal $G$-bundle away from the nodes,
  • 4)
    if the node $z$ lies over the node $x$ of the corresponding fibre of $\mathcal{X}\rightarrow S$, and $z_{1}$ and $z_{2}$ are local coordinates of the two branches near $z$, then the generator $h$ of the stabilizer $Stab_{G}(z)$ acts as
    $h.(z_{1},z_{2})=(\zeta z_{1},\zeta^{-1}z_{2}),$
    where $d=\operatorname{ord}(h)$ and $\zeta$ is a primitive $d$-th root of the unity.
When $S=\operatorname{Spec}(\mathbb{C})$, we say that $f:Z\rightarrow X$ is an admissible $G$-cover. in this case, let $\tilde{f}:\tilde{Z}\rightarrow\tilde{X}$ be the lifting of $f$ on the normalisations. Then for each node of $Z$, condition $(4)$ is a compatibility condition for the local monodromies of $\tilde{f}$ at the two points in $\tilde{Z}$ lying above the node in $Z$.
Now, let $G$ be a cyclic group of order $d$ with generator $\sigma$. In the following, let
$f_{D}:\widetilde{D}\longrightarrow D,$
be the admissible $G$-cover satisfying the following:
  • the curves $\widetilde{D}$ and $D$ are connected stable curves of genus $g$ and $g^{\prime}$, respectively;
  • the curve $\widetilde{D}$ has exactly one node, and this node is fixed by the action of $G$;
  • the normalization of $f_{D}$ is the $G$-cover
    $f_{D}^{\nu}:\widetilde{D}^{\nu}\longrightarrow D^{\nu},$
    with monodromy data $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{r}).$
More concretely, let $p_{1},p_{2}\in D^{\nu}$ be the two preimages of the node of $D$. These points are the branch points of the normalized cover $f_{D}^{\nu}$, with local monodromies $1$ and 1-1, respectively. The admissible cover $f_{D}$ is obtained from its normalization $f_{D}^{\nu}$ by gluing the points $p_{1}$ and $p_{2}$ in $D^{\nu}$ to form the node of $D$, and simultaneously gluing their corresponding preimages in $\tilde{D}\nu$. Note that we have $g(\tilde{D}^{\nu})=g-1$ and $g(D^{\nu})=g^{\prime}-1$.
We now study the Prym variety associated with the covering $f_{D}$, inspired by [2, 9, 18].
At the node $s$ of $\tilde{D}$, we make the usual identification:
$\mathcal{K}^{*}_{s}/\mathcal{O}_{s}^{*}\cong\mathbb{C}^{*}\times{\mathbb{Z}}\times{\mathbb{Z}},$
and the induced action by $\sigma$ works as:
$\sigma^{*}((z,m,n)_{s})=(\rho^{(m+n)}z,m,n)_{s},$
where $\rho$ is a primitive $d$-th root of the unity. We start with the following:
Lemma 3.2.
[…]
$L\simeq M\otimes\sigma^{*}M^{-1},$
[…]
Proof.
We need to describe divisors $H$ such that $L\simeq\mathcal{O}_{\tilde{D}}(H)$ and $(f_{D})_{*}(H)=0$. Writing $H=\sum_{x\in\tilde{D}_{reg}}x+(z,m,n)_{s}$, we get that $H$ must be a linear combination of two kinds of contributions: from smooth points $x\in\tilde{D}_{\mathrm{reg}}$, divisors of the form $x-\sigma^{j}x$ ($j\in[0,\ldots,d-1]$), and at the singular point $s\in\tilde{D}$, divisors of the form $(\rho^{k},0,0)_{s}$. Indeed, using that
$(f_{D})_{*}((z,m,n)_{s})=(\Pi_{i=0}^{d-1}(\sigma^{i})^{*}z,m,n)_{s}=\begin{cases}(z^{d},m,n)_{f(s)}\quad\quad\quad\text{if $d$ odd}\\ ((-1)^{(m+n)}z^{d},m,n)_{f(s)}\quad\text{if $d$ even}\end{cases}$
we have that
$(f_{D})_{*}((\rho^{k},0,0)_{s})=(1,0,0)_{f(s)}.$
Since $(\rho^{k},0,0)_{s}=\sum(\rho,0,0)_{s}$, it suffices to show that $(\rho,0,0)_{s}$ is in the image of $1-\sigma^{*}.$ This follows from
$(1,-1,0)-\sigma^{*}(1,-1,0)=(\rho,0,0).$
Hence $H=E-\sigma^{*}E,$ for some divisor $E\in\tilde{D}$.
Let $P$ denote the kernel of the norm map $\operatorname{Nm}:J\tilde{D}\rightarrow JD$. Lemma 3.2 implies that $P$ is the variety of line bundles in $\ker\operatorname{Nm}$ of the form $M\otimes\sigma^{*}M^{-1}$. The Prym variety $P(\tilde{D},D)$ of $f_{D}$ is, by definition, the connected component of the identity of $P$. We have the following:
Theorem 3.3.
The Prym variety $P(\tilde{D},D)$ is isogenous to the Prym variety $P(\tilde{D}^{\nu},D^{\nu})$ associated with the normalized covering $\tilde{D}^{\nu}\rightarrow D^{\nu}$.
Proof.
We consider the following diagram of commutative algebraic groups:
[…]
where $\tilde{T},T$ are the groups of classes of divisors of degree 0 and singular support. Since $(f_{D})^{*}$ is injective on $T$ and $\operatorname{Nm}\circ(f_{D})^{*}=d$, the norm on $\tilde{T}$ is surjective and
$\ker\operatorname{Nm}|_{\tilde{T}}\cong\tilde{T}_{d}:=\{\text{points in $\tilde{T}$ whose order divides $d$}\}.$
Hence, taking the kernels in diagram (3.1), we obtain the exact sequence:
(3.2) $0\rightarrow\tilde{T}_{d}\rightarrow P\rightarrow R\rightarrow 0,$
where $R:=\ker\operatorname{Nm}J\tilde{D}^{\nu}\rightarrow JD^{\nu}$. Since $\tilde{D}^{\nu}\rightarrow D^{\nu}$ is ramified, $R$ is an abelian variety, and indeed, it is the Prym variety $P(\tilde{D}^{\nu},D^{\nu}).$ Moreover, since $\dim\tilde{T}=\dim T=1$, we have that $\tilde{T}_{d}\cong{\mathbb{Z}}/d$. Therefore, using (3.2), we get that $P$ and $R$ are isogenous. ∎
4. Maximal monodromy
In this section we prove a criterion for the maximality of the algebraic monodromy group for certain families of cyclic coverings. We start by describing the natural $\mathbb{Q}$-variation of Hodge structures associated with a family of Galois coverings.
4.1.
Let $G$ be a finite cyclic group. As explained in Section 2, let us consider a topological type of $G$-action and let $f:\mathcal{C}\rightarrow B$ be the corresponding algebraic family of curves parametrizing all Galois covers carrying a $G$-action topologically equivalent to the one fixed. Thus every cover $C_{b}\to C^{\prime}_{b}=C_{b}/G$ has the prescribed invariants: the genera $g=g(C_{b})$ and $g^{\prime}=g(C^{\prime}_{b})$, together with the chosen ramification and monodromy data. Let
$\mathbb{V}=R^{1}f_{*}{\mathbb{Q}}$
be the local system underlying the natural $\mathbb{Q}$}-variation of Hodge structure (VHS) associated with this family. Fix a Hodge-generic point $b\in B$ with respect to the variation $\mathbb{V}$. The $G$-action on the family induces a compatible action on the VHS. Indeed, denote by $X(G)$ the set of $\mathbb{C}$-irreducible characters of $G$. Then the complexified local system decomposes as
$\mathbb{V}_{\mathbb{C}}=\bigoplus_{\chi\in X(G)}\mathbb{V}_{\chi},$
where $\mathbb{V}_{\chi}$ is the $\chi$-isotypical summand. At the fiber $b$, we have $\mathbb{V}_{b}=H^{1}(C_{b},{\mathbb{Q}})$, and after extension of scalars to $\mathbb{C}$ we obtain
$H^{1}(C_{b},\mathbb{C})=H^{1,0}(C_{b})\oplus H^{0,1}(C_{b}).$
Moreover, the space of holomorphic differentials admits a decomposition
(4.1) $H^{1,0}(C_{b})=\bigoplus_{\chi\in X(G)}m_{\chi}V_{\chi},$
where $V_{\chi}$ denotes the irreducible $\mathbb{C}$-representation corresponding to $\chi$ and $m_{\chi}$ its multiplicity. In particular,
$\mathbb{V}_{\chi,b}=H^{1}(C_{b},\mathbb{C})_{\chi}=m_{\chi}V_{\chi}\;\oplus\;\overline{m_{\chi}V_{\chi}}=m_{\chi}V_{\chi}\oplus m_{\bar{\chi}}V_{\bar{\chi}}.$
Let $X(G,\mathbb{R})$ be the set of $\mathbb{R}$-irreducible representations of $G$. For a character $\chi\in X(G)$, we denote by $(\chi,\overline{\chi})$ the corresponding real representation. Then one has a decomposition over $\mathbb{R}$ of the form
$\mathbb{V}_{\mathbb{R}}=\bigoplus_{(\chi,\overline{\chi})\in X(G,\mathbb{R})}\mathbb{V}_{(\chi,\overline{\chi})},$
and if $\chi\neq\overline{\chi}$, we have
$\mathbb{V}_{(\chi,\overline{\chi})}\otimes_{\mathbb{R}}\mathbb{C}=\mathbb{V}_{\chi}\oplus\mathbb{V}_{\overline{\chi}},$
whereas if $\chi=\overline{\chi}$ one obtains
$\mathbb{V}_{(\chi,\chi)}\otimes_{\mathbb{R}}\mathbb{C}=\mathbb{V}_{\chi}.$
The polarization on the VHS restricts as follows: when $\chi\neq\overline{\chi}$, it induces a Hermitian form of signature $(m_{\chi},m_{\overline{\chi}})$ on $\mathbb{V}_{(\chi,\overline{\chi})}$; in the case $\chi=\overline{\chi}$, it yields a symplectic form on $\mathbb{V}_{(\chi,\chi)}$ (see [7, Corollary 2.21–2.23]).
Let $\operatorname{Mon}^{0}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ denote the connected monodromy group, and $\operatorname{Hg}\subset\operatorname{GL}(H^{1}(C_{b},{\mathbb{Q}}))$ the generic Hodge group of the VHS. By a theorem of André [1, Theorem 1], $\operatorname{Mon}^{0}$ is a semisimple normal subgroup of $\operatorname{Hg}$. More precisely, we have
(4.2) $\operatorname{Mon}^{0}\;\subseteq\;\operatorname{Hg}^{der}\;\subseteq\;\ (\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der},$
where $\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G}$ is the centralizer of $G$ inside $\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))$. Upon extension of scalars to $\mathbb{R}$, the latter admits the following explicit decompositions:
(4.3) $\operatorname{Sp}(H^{1}(C_{b},\mathbb{R}))^{G}\simeq\prod_{(\chi,\bar{\chi})\in X(G,\mathbb{R})}\operatorname{Sp}(H^{1}(C_{b},\mathbb{R})_{\chi})^{G}\simeq\prod_{(\chi,\bar{\chi}),\chi\neq\bar{\chi}}\operatorname{U}(m_{\chi},m_{\bar{\chi}})\times\prod_{(\chi,\bar{\chi}),\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
We denote by $\operatorname{Mon}^{0}(\chi)$ the projection of $\operatorname{Mon}^{0}_{\mathbb{R}}$ onto the $\chi$-factor of the decomposition. If $\chi\neq\bar{\chi}$, we have $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$, whether for $\chi=\bar{\chi}$, then $\operatorname{Mon}^{0}(\chi)\subseteq\operatorname{Sp}_{2m_{\chi}}.$ Thus, we have
(4.4) $\operatorname{Mon}^{0}_{\mathbb{R}}\hookrightarrow\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
4.2.
For the remaining part of this section, we focus on the inclusion (4.4) and we prove a criterion for the maximality of the monodromy. We will work with families satisfying the following condition:
Definition 4.3.
The family has no repeating factors in the monodromy if for all $\chi_{i},\chi_{j}\in X(G),i\neq j$, different from the trivial representation and with $m_{\chi_{i}}m_{\bar{\chi}_{i}}m_{\chi_{j}}m_{\bar{\chi}_{j}}\neq 0$, we have that:
  • $\operatorname{SU}(m_{\chi_{i}},m_{\bar{\chi}_{i}})\not\cong\operatorname{SU}(m_{\chi_{j}},m_{\bar{\chi}_{j}})$ if $\chi_{i}\neq\bar{\chi}_{i}$ and $\chi_{j}\not\cong\bar{\chi}_{j}$
  • $\operatorname{Sp}(2m_{\chi_{i}},\mathbb{R})\neq\operatorname{Sp}(2m_{\chi_{j}},\mathbb{R})$ if $\chi_{i}=\bar{\chi}_{i}$ and $\chi_{j}=\bar{\chi}_{j}$.
In other words, a family has no repeating factors in the monodromy if all the factors appearing in $(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}$ are mutually distinct. Our main theorem uses an induction argument on the genus $g^{\prime}$ of the base curve $C^{\prime}_{b}$. The base of the induction, i.e. the case with $g^{\prime}=0$, is the following result:
Theorem 4.4.
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings of the line with no repeating factors in the monodromy. Then the monodromy is maximal, that is
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad}.$
This result is proved in [26, Theorem 1.7]. A fundamental ingredient, due to Rohde [24], is that for cyclic coverings of the line, whenever the multiplicities $m_{\chi}$ and $m_{\bar{\chi}}$ are nonzero, one has $\operatorname{Mon}^{0}(\chi)=\operatorname{SU}(m_{\chi},m_{\bar{\chi}})$ if $\chi\neq\bar{\chi}$, and $\operatorname{Mon}^{0}(\chi)=\operatorname{Sp}_{2m_{\chi}}$ if $\chi=\bar{\chi}$.
The induction step in our proof is based on a degeneration argument. More precisely, we consider admissible $G$-coverings that sit in the boundary of the family $f:\mathcal{C}\rightarrow\mathsf{B}$. The general element in the family is a covering $C\rightarrow C^{\prime}$ with $g(C)=g$, $g(C^{\prime})=g^{\prime}$ and monodromy $\Theta=(\theta_{1},\ldots,\theta_{s}).$ We degenerate this covering to an admissible $G$-cover $f_{D}:\tilde{D}\rightarrow D$ of a connected, stable, 1-nodal curve $D$ of arithmetic genus $g^{\prime}$, whose covering $\tilde{D}$ has exactly one node on which is fixed by $G$ (and mapped to the node of $D$). The normalization of $f_{D}$ is the covering $\tilde{D}^{\nu}\rightarrow D^{\nu}$ with monodromy $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{s})$. This is the setup of section 3.
Remark 4.5.
Note that the covering $\tilde{D}\rightarrow D$ is the generic element of an irreducible component of the intersection $\overline{\mathsf{M}}\cap\Delta_{0},$ where $\mathsf{M}$ is the moduli image of the family $f$ in $\mathcal{M}_{g}$, $\overline{\mathsf{M}}$ its closure in $\overline{\mathcal{M}}_{g}$, and $\Delta_{0}$ the boundary divisor of irreducible nodal curves. We will refer to the family of the coverings $\tilde{D}\rightarrow D$ as the $\Delta_{0}$-degeneration of $f$.
As in (4.1), let $m_{\chi}$ be the multiplicity of the $\mathbb{C}$-irreducible representation corresponding to the character $\chi$ of the group $G$ acting on $H^{1,0}(C)$. Similarly, let $m_{\chi}^{\nu}$ be the dimension of the $\chi$-eigenspace for the action of $G$ on $H^{1,0}(\tilde{D}^{\nu})$. The Lemma below compares the dimensions of the eigenspaces.
Lemma 4.6.
The following equality holds
$m_{\chi_{i}}=m_{\chi_{i}}^{\nu}$
for all $\chi\in X(G)$ different from the trivial representation.
Proof.
The result follows directly from the Chevalley-Weil formula (2.1). Indeed, using that $\Theta^{\nu}$ is defined with the same $s$-tuple as $\Theta$ plus the pair (1,1)(-1,1), we have
$m^{\nu}_{\chi_{i}}=-1+m_{\chi_{i}}+1.$
We are now ready to state our main theorem.
Theorem 4.7.
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then the monodromy is maximal, that is
(4.5) $\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der,ad}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{ad},$
and, in particular, if $g^{\prime}\geq 2$
(4.6) $\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C_{b},{\mathbb{Q}}))^{G})^{der}.$
Proof.
Let $C_{b}\rightarrow C^{\prime}_{b}=C_{b}/G$ be the generic element in the family. We use induction on the genus $g^{\prime}$ of the base curve. For $g^{\prime}=0$, the statement follows by Theorem 4.4. Let us show that the statement holds for $g^{\prime}>0$, assuming that it is true for $g^{\prime}-1$. Since $JC_{b}\sim JC_{b}^{\prime}\times\operatorname{Prym}(C_{b},C_{b}^{\prime})$, we have
$\operatorname{Mon}^{0}(JC_{b})\subseteq\operatorname{Mon}^{0}(JC^{\prime}_{b})\times\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))\subseteq\operatorname{Sp}_{2g^{\prime}}\times\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}),$
where the first term in the product comes from the trivial representation. As $C_{b}^{\prime}$ moves in the whole moduli space $\mathcal{M}_{g^{\prime}}$, we have $\operatorname{Mon}^{0}(JC^{\prime}_{b})=\operatorname{Sp}_{2g^{\prime}}$. Hence, to prove maximality it is enough to focus on $\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$. We first observe that we have
$\operatorname{Mon}^{0,ad}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))\subseteq\prod_{\chi\neq\bar{\chi},\\ \ m_{\chi}m_{\bar{\chi}}\neq 0}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi},\\ m_{\chi}\neq 0}\operatorname{Sp}(2m_{\chi},\mathbb{R}),$
and using a degeneration, we can show that we actually have equality. Indeed, as in 4.2, we degenerate to the admissible cover $\tilde{D}\rightarrow D$ whose normalization $\tilde{D}^{\nu}\rightarrow D^{\nu}$ is endowed with monodromy $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{s})$. By Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. By Lemma 4.6, the covering $\tilde{D}^{\nu}\rightarrow D^{\nu}$ has
$\operatorname{SU}(m^{\nu}_{\chi_{i}},m^{\nu}_{\bar{\chi}_{i}})\neq\operatorname{SU}(m^{\nu}_{\chi_{j}},m^{\nu}_{\bar{\chi}_{j}})$
for all $\chi_{i},\chi_{j}\in X(G)$ different from the trivial representation, $i\neq j$ with $m^{\nu}_{\chi_{i}}m^{\nu}_{\bar{\chi}_{i}}\neq 0$. In other words, $\tilde{D}^{\nu}\rightarrow D^{\nu}$ also has no repeating factors in the monodromy. Since $g(D^{\nu})=g^{\prime}-1$, the induction hypothesis implies that $\operatorname{Mon}^{0}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu}))$ is maximal. Moreover we have $\operatorname{Mon}^{0,ad}(\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu}))=\operatorname{Mon}^{0,ad}(\operatorname{Prym}(\tilde{D},D))\subseteq\operatorname{Mon}^{0,ad}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$, which then gives maximality for $\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))$, i.e. (4.5). Using Chevalley-Weil formula 2.1, we get that no 0-dimensional eigenspace appears if $g^{\prime}\geq 2$. Thus, (4.6) follows. ∎
5. The Coleman-Oort conjecture
In this section, we discuss some consequences of our main result in relation to the Coleman-Oort conjecture.
5.1.
As recalled in the introduction, for low genus, there exist examples of positive dimensional special subvarieties $S\subset\mathcal{A}_{g}$ that lie generically in the Torelli locus. The main construction proceeds as follows. As in section 2, let $G$ be a finite group, and consider the family $\mathcal{C}\to\mathsf{B}$ of all Galois covers with a fixed topological type, set $\mathsf{M}\subset\mathcal{M}_{g}$ to be the image of the base $\mathsf{B}$ under the moduli map, and $\mathsf{Z}:=\overline{j(\mathsf{M})}\subset\mathcal{A}_{g}$ as the closure of the image under the Torelli map.
Let $S(G)\subset\mathcal{A}_{g}$ denote the PEL-type special subvariety associated with the group $G$. This is the special subvariety whose uniformizing symmetric space corresponds to the group $(\operatorname{Sp}_{2g})^{G}$. By construction, we have the inclusion $\mathsf{Z}\subseteq S(G)$. It is shown in [22] for the case $g^{\prime}=0$ and $G$ cyclic, and extended in [10, 11] to arbitrary $G$ and $g^{\prime}$, that if the following numerical condition holds for a general $C]\in\mathsf{M}$:
($\star$) $\dim\mathsf{Z}=\dim H^{0}(C,2K_{C})^{G}=\dim\left(S^{2}H^{0}(C,K_{C})\right)^{G}=\dim S(G),$
then $\mathsf{Z}$ is a special subvariety of $\mathcal{A}_{g}$. Since $\mathsf{Z}$ lies in the Torelli locus, this gives a counterexample to the Coleman-Oort conjecture whenever ($\star$5.1) is satisfied.
5.2.
It is important to note that the condition ($\star$5.1) is a priori only sufficient: if $\dim\mathsf{Z}<\dim S(G)$, one cannot conclude that $\mathsf{Z}$ is not special. Indeed, if we denote by $S_{f}$ be the smallest special subvariety of $\mathcal{A}_{g}$ containing $\mathsf{Z}$, we have
$\mathsf{Z}\subseteq S_{f}\subseteq S(G),$
and, by definition, $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}$. Proving the necessity of condition ($\star$5.1) would thus amount to showing that $\mathsf{Z}$ is special if and only if $\mathsf{Z}=S_{f}=S(G)$.
5.3.
To complete the picture, we recall that special subvarieties are subvarieties in $\mathcal{A}_{g}$, which are linear with respect to the locally symmetric metric induced from the Siegel space $\mathfrak{S}_{g}$. More precisely, denote by $\pi:\mathfrak{S}_{g}\rightarrow A_{g}$ the natural projection map and assume that an algebraic subvariety $Z\subset\mathcal{A}_{g}$ is totally geodesic if $Z=\pi(X)$, for a totally geodesic submanifold $X\subset\mathfrak{S}_{g}$. Then, special subvarieties are totally geodesic (and not all totally geodesic subvariety of $\mathcal{A}_{g}$ are special). Thus, if we denote by $X_{f}$ the smallest totally geodesic subvariety of $\mathcal{A}_{g}$ containing our family $\mathsf{Z}$, we have:
$\mathsf{Z}\subseteq X_{f}\subseteq S_{f}\subseteq S(G).$
We recall that $X_{f}$ is the totally geodesic subvariety whose uniformizer is the symmetric space associated with the real group $\operatorname{Mon}^{0}_{\mathbb{R}}$, while $S_{f}$ is the special subvariety associated with the generic Hodge group $\operatorname{Hg}_{\mathbb{R}}$ of the family (see [21]). Hence, the inclusions above reflect the inclusions in (4.2).
In this setting, our Theorem 4.7, reads as follows.
Proposition 5.4.
Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. Then $X_{f}=S_{f}=S(G)$. In particular, $\mathsf{Z}$ is totally geodesic if and only if it is special and if and only if ($\star$5.1) holds.
This result allows us to go further in the classification of the counterexamples to the Coleman-Oort conjecture. Indeed in [5], employing a computational approach complemented by theoretical arguments, the authors proved that the positive-dimensional families of Galois covers satisfying ($\star$5.1) with $2\leq g\leq 100$ are only those listed in [10, 11]. Our next result is an improvement of the classification in the case of families of cyclic coverings with no repeating factors in the monodromy. Indeed, we can prove the following:
Theorem 5.5.
The only positive dimensional totally geodesic families of cyclic coverings with no repeating factors in the monodromy are the ones appearing in [10, 11]. In particular, for $g\geq 8$ there are no totally geodesic families of this kind.
In order to prove this result, we first need a preliminary result concerning families of cyclic coverings of genus 1 curves. Our notation will be as in Remark 4.5, namely $f:\mathcal{C}\rightarrow\mathsf{B}$ is a family of cyclic coverings with monodromy $\Theta=(\theta_{1},\dots,\theta_{s})$, $\tilde{D}\rightarrow D$ is the family of $\Delta_{0}$-degeneration of $f$ and $\tilde{D}^{\nu}\rightarrow D^{\nu}$ is the family of their normalisations. This is a family $f^{\nu}:\mathcal{D}^{\nu}\rightarrow\mathsf{B}^{\nu}$ with monodromy $\Theta^{\nu}=(1,-1,\theta_{1},\dots,\theta_{s})$.
Proposition 5.6.
Let $f:\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic coverings of genus-$1$ curves with no repeating factors in the monodromy. Suppose that $f$ satisfies ($\star$5.1), then $f^{\nu}:\mathcal{D}^{\nu}\rightarrow\mathsf{B}^{\nu}$ satisfies it as well.
Proof.
First observe that $\dim\mathsf{B}^{\nu}=\dim\mathsf{B}-1$. We denote by $S^{\nu}(G)$ the PEL special subvariety associated with the family $f^{\nu}$. To conclude we need to show that $\dim S^{\nu}(G)=\dim S(G)-1$. We write $JC_{b}\sim JC_{b}^{\prime}\times\operatorname{Prym}(C_{b},C_{b}^{\prime})$. Since ($\star$5.1) holds for the family $f$, we have
(5.1) $\operatorname{Mon}^{0}(JC_{b})=\operatorname{Mon}^{0}(JC^{\prime}_{b})\times\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\operatorname{Sp}_{2}\times\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
In particular,
$\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
By Lemma 4.6, the covering $D^{\nu}\rightarrow D^{\nu}$ has still no repeating factions in the monodromy. Furthermore, by Theorem 3.3, we have $\operatorname{Prym}(\tilde{D},D)\sim\operatorname{Prym}(\tilde{D}^{\nu},D^{\nu})$. Also, recall that $g(D^{\nu})=0$ and hence $JD^{\nu}\sim\operatorname{Prym}(\tilde{D},D)$. Therefore, by Theorem 4.4
$\operatorname{Mon}^{0}(JD^{\nu})=\operatorname{Mon}^{0}(\operatorname{Prym}(C_{b},C_{b}^{\prime}))=\prod_{\chi\neq\bar{\chi}}\operatorname{SU}(m_{\chi},m_{\bar{\chi}})\times\prod_{\chi=\bar{\chi}}\operatorname{Sp}(2m_{\chi},\mathbb{R}).$
Comparing with (5.1) and taking the orbit spaces, we get that $\dim S^{\nu}(G)=\dim S(G)-1$. ∎
Proof of Theorem 5.5 Let $\mathcal{C}\rightarrow\mathsf{B}$ be a family of cyclic $G$-coverings with no repeating factors in the monodromy. By Proposition 5.4, the family is totally geodesic if and only if it is special and if and only if it satisfies condition ($\star$5.1). In [13], it is shown that there are no special families satisfying ($\star$5.1) for $g^{\prime}\geq 2$. On the other hand, the classification of such families with $g^{\prime}=0$ follows from the main theorem in [22]. All these examples are in genus $g\leq 7$. Hence, let us focus on the case $g^{\prime}=1$. In [11], 4 examples are provided. Again, they all have $g\leq 7$. Therefore, we only have to prove that no other further examples exist. Suppose that such a family $f$ exists. By Proposition 5.6, if follows that the normalization $f^{\nu}$ still satisfies ($\star$5.1) and hence it is one of the families listed in [22]. Since $g(\tilde{D}^{\nu})=g(C)-1$, we obtain $g\leq 8$. Therefore, the statement follows from the main theorem in [5], where the authors prove that the only positive-dimensional families of Galois covers satisfying ($\star$5.1) with $2\leq g\leq 100$ are only those listed in [10, 11]. ∎
Remark 5.7.
Notice that, thanks to Proposition 5.4, we may work with totally geodesic subvarieties rather than only with special ones. In other words, our results contribute to the stronger form of the Coleman–Oort conjecture, which predicts the non-existence of positive-dimensional totally geodesic subvarieties of $\mathcal{A}_{g}$ that are generically contained in the Torelli locus.
We now show that, dealing with special families, our assumption on the monodromy is not too strong. Indeed, we have the following:
Theorem 5.8.
All known examples of special families of cyclic coverings have no repeating factors in the monodromy.
Proof.
The examples of special families of cyclic coverings can be found in [10, 11]. The proof of the statement is based in the following considerations. First, if the family is one-dimensional, then the uniformizing symmetric space of $S(G)$ is $\mathfrak{S}_{1}$, so this is automatic. Families with double and triple coverings always have no repeating factors in the monodromy; this is discussed in section 6. Finally, for the remaining higher dimensional families we refer the reader to [27, Theorem 1.2], where it is shown that their uniformizing symmetric space is an irreducible symmetric space, and this implies that only one factor appears in $(\operatorname{Sp}^{G})^{ad}$ (hence, in particular, they have no repeating factors in the monodromy). ∎
5.9.
It is interesting to compare our results with Theorem 1.3 in [17]. Indeed, in this paper, under some numerical assumptions, the authors show that the identity component of the Zariski closure of the image of the monodromy map associated with a family of $G$-covers is the is the commutator subgroup of $(\operatorname{Sp}_{2g})^{G}$. More precisely, in the case of a family of abelian covers $C\rightarrow C^{\prime}$ the result provides
$\operatorname{Mon}^{0}=(\operatorname{Sp}(H^{1}(C,{\mathbb{Q}}))^{G})^{der}$
as soon as $g(C^{\prime})\geq 4$. This implies
$\operatorname{Mon}^{0,ad}=(\operatorname{Sp}(H^{1}(C,{\mathbb{Q}}))^{G})^{ad},$
namely it guarantees the maximality of the monodromy in the sense of our Theorem 4.7. Hence, for $g^{\prime}\geq 4$ our result provides only an alternative technique to get the maximality in our hypothesis. On the other hand, our result also covers the genera $g^{\prime}\leq 4$ for cyclic coverings with no repeating factors in the monodromy. We also note that [17, Theorem 1.3] implies that $S_{f}=S(G)$ for every family of abelian covers with $g^{\prime}\geq 4$. Thus, arguing as in the proof of Theorem 5.5 and, using that by [13, Theorem 1.1] if $g^{\prime}\geq 2$ there are no families satisfying condition ($\star$5.1), then we can exclude the existence of families of abelian covers with $g^{\prime}\geq 4$ yielding special (or even totally geodesic) subvarieties generically contained in the Torelli locus.
6. Application to double and triple coverings
In this Section we apply Proposition 5.4 and Theorem 5.5 to double and triple coverings of curves of a fixed genus. We start by fixing some notation. Let $\mathcal{B}_{g,h}$ be the image in $\mathcal{M}_{g}$ of the family of all double covers $f:C\rightarrow C^{\prime}$ with $g(C)=g$ and $g(C^{\prime})=h$. Notice that, by Riemann-Hurwitz, the number of ramification points is r=2g4h+2r=2g-4h+2. It follows that $\operatorname{dim}(\mathcal{B}_{g,h})=2g-h-1$.
Corollary 6.1.
The smallest totally geodesic subvariety containing $\mathcal{B}_{g,h}$ is the PEL special variety $S(\mathcal{B}_{g,h})\cong\mathcal{A}_{h}\times\mathcal{A}_{g-h}.$ In particular, $\overline{j(\mathcal{B}_{g,h})}$ is not totally geodesic in $\mathcal{A}_{g}$ whenever $(g,h)\neq(1,0),(2,0),(2,1),(3,1)$.
Proof.
We start by observing that double coverings families obviously have no repeating factors in the monodromy, since the group $G=\mathbb{Z}/2\mathbb{Z}$ has only one non-trivial representation. Hence Proposition 5.4 implies that the smallest special subvariety containg the moduli image of the family is the PEL $S(G)$. To compute it, we use the Chevalley-Weil formula and obtain that
$\operatorname{Mon}^{0}(\mathcal{B}_{g,h})=\operatorname{Hg}_{\mathbb{R}}(\mathcal{B}_{g,h})=\operatorname{Hg}_{\mathbb{R}}(S(\mathcal{B}_{g,h}))=\operatorname{Sp}_{2h}\times\operatorname{Sp}_{2(g-h)}.$
Thus, $\mathcal{B}_{g,h}$ is totally geodesic if and only if $\dim\mathcal{B}_{g,h}=\dim\mathcal{A}_{h}\times\mathcal{A}_{g-h}$; namely if and only if $(g,h)=(1,0),(2,0),(2,1),(3,1)$. ∎
Now let $\mathcal{T}_{g,h}(r,m)$ be the image in $\mathcal{M}_{g}$ of the family of all Galois triple coverings $f:C\rightarrow C^{\prime}$ with $g(C)=g,\ g(C^{\prime})=h$, $r=2m+n$ ramification points and fixed monodromy datum, which we write as:
(6.1) $\Theta=(x_{1},\ldots,x_{2m},y_{1},\ldots,y_{n}),$
where $x_{i}+x_{i+1}=0$ mod $3$ and $y_{i}+y_{j}\neq 0$ mod $3$ for all $i$ and $j$. A couple $(x_{i},x_{i+1})$ with $x_{i}+x_{i+1}=0$ mod $3$ is called a canceling pair. With the notation in (6.1), $m$ is the number of the canceling pairs in $\Theta$, and $\sum_{i}y_{i}=0\mod 3$.
We observe that (since it is a triple covering), the choice of $r$ and $m$ uniquely determines $\Theta$. In particular, we will assume that $(y_{1},\ldots,y_{n})=([1]_{3},\ldots,[1]_{3})$, and thus $r=2m\mod 3$. Recall that, by Riemann-Hurwitz, we have r=g3h+2r=g-3h+2. We first consider the unramified case, i.e. we assume $r=0,h\geq 2$. We will denote by $\mathcal{A}(p,q)$ a totally geodesic subvariety of $\mathcal{A}_{g}$ whose uniformizing symmetric is associated with the group $SU(p,q)$. We have the following:
Corollary 6.2.
The smallest totally geodesic subvariety containing $\mathcal{T}_{3h-2,h}$ is the PEL special variety $S(\mathcal{T}_{3h-2,h})\cong\mathcal{A}_{h}\times\mathcal{A}(h-1,h-1)$. In particular, $\overline{j(\mathcal{T}_{3h-2,h})}$ is not totally geodesic in $\mathcal{A}_{g}$.
Proof.
As before, we observe that triple coverings families have no repeating factors in the monodromy, as the group $G=\mathbb{Z}/3\mathbb{Z}$, has only one pair of non-trivial representations $(\chi_{[1]},\chi_{[2]})$, which are conjugate to each other. Thus, by Proposition (5.4), the smallest special subvariety containing $\mathcal{T}_{3h-2,h}$ is $S(G)$. We use the Chevalley-Weil formula and obtain that
$\operatorname{Mon}^{0}(\mathcal{T}_{3h-2,h})=\operatorname{Hg}_{\mathbb{R}}(\mathcal{T}_{3h-2,h})=\operatorname{Hg}_{\mathbb{R}}(S(\mathcal{T}_{3h-2,h}))=\operatorname{Sp}_{2h}\times\,SU(h-1,h-1),$
which shows that $S(\mathcal{T}_{3h-2,h})$ is as claimed. By dimension reasons, we get that $\mathcal{T}_{3h-2,h}\subsetneq S(\mathcal{T}_{3h-2,h})$, and thus is not totally geodesic. ∎
Finally, we discuss triple Galois ramified coverings. We have the following:
Corollary 6.3.
The smallest totally geodesic subvariety containing $\mathcal{T}_{g,h}(r,m)$ is the PEL special variety $S(\mathcal{T}_{g,h}(r,m))\cong\mathcal{A}_{h}\times\mathcal{A}(d_{1},d_{2})$, where $d_{1}=h-1+\frac{2}{3}r-\frac{1}{3}m$ and $d_{2}=h-1+\frac{1}{3}r+\frac{1}{3}m$. Hence, $\overline{j(\mathcal{T}_{g,h}(r,m))}$ is totally geodesic in $\mathcal{A}_{g}$ if and only if
(g,h,r,m)=(2,0,4,2),(3,0,5,1),(4,0,6,0),(3,1,2,1),(4,1,3,0).(g,h,r,m)=(2,0,4,2),(3,0,5,1),(4,0,6,0),(3,1,2,1),(4,1,3,0).
Proof.
As before, we use Proposition 5.4 and the Chevalley-Weil formula and obtain that
$\operatorname{Mon}^{0}(\mathcal{T}_{g,h}(r,m))=\operatorname{Hg}_{\mathbb{R}}(\mathcal{T}_{g,h}(r,m))=\operatorname{Sp}_{2h}\times SU(d_{1},d_{2}),$
with $d_{1}$ and $d_{2}$ as in the statement, which shows that $S(\mathcal{T}_{g,h})$ is as claimed. Since the smallest totally geodesic subvariety containing $\mathcal{T}_{g,h}(r,m)$ is the PEL, we have that the family is totally geodesic if and only if $\dim\mathcal{T}_{g,h}=\dim S(\mathcal{T}_{g,h})$. When $h\geq 2$ this never happens by [13], while for $h=0,1$, this is satisfied if and only if it is one of the families in [10, 12], namely it is one of the family listed in the statement. ∎